OpenStudy (anonymous):

What type of line is the graph of 4x + y = -3?

OpenStudy (anonymous):

like a positive or negative line? or what do you mean?

OpenStudy (anonymous):

what are the choices?

OpenStudy (anonymous):

I used the values 1, 0, and -1. 4(1) + y = -3 = -7 (1, -7) 4(0) + y = -3 = -3 (0, -3) 4(-1) +y = -3 = 1 (-1, 1) My options: Vertical Horizontal xRising Falling The one with the x is my choice.

OpenStudy (anonymous):

No, wait, it's falling. It's a negative slope.

OpenStudy (anonymous):

you dont need to plug in anything simply subtract 4x from both sides and yo jat slope intercept,